If P is a fraction approximated thy 0.18 with relative error of 0.04 find P
1
Expert's answer
2020-09-21T14:46:16-0400

"Rel(P)=\frac{P-x}{P}"

"0.04=\frac{P-0.18}{P}"

"0.04P=P-0.18"

"P=\frac{0.18}{1-0.04}=\frac{3}{16}=0.1875"